OpenStudy (anonymous):

Wendy can wash a car twice as fast as Jason. When they work together, Wendy and Jason can wash a large van in 2 hours. How many hours would it take Jason to wash the van by himself?

OpenStudy (anonymous):

j rate of jason (j + 2j)2 = 1 j = 1/6 so 1/(1/6) = 6 hrs

OpenStudy (anonymous):

the choices are 5,4,3,2.. @sourwing

OpenStudy (anonymous):

then it doesn't contain 6hrs

OpenStudy (anonymous):

huh? @sourwing

OpenStudy (anonymous):

@***[ISURU]*** can you help??